# Pseudo-Physics-Informed Neural Operators: Enhancing Operator Learning from Limited Data

### v1

@@ 4 Related Work
Operator learning is a rapidly evolving research field, with a variety of methods categorized as neural operators. Alongside FNO, several notable approaches have been introduced, such as low-rank neural operator (LNO) (Li et al., 2020b, ), multipole graph neural operator (MGNO) (Li et al., 2020a, ), multiwavelet-based NO (Gupta et al.,, 2021), and convolutional NOs (CNO) (Raonic et al.,, 2023). Deep Operator Net (DON) (Lu et al.,, 2021) is another popular approach, consisting of a branch network applied to input function values and a trunk network applied to sampling locations. The final prediction is obtained through the dot product of the outputs from the two networks. To enhance stability and efficiency, Lu et al., (2022) proposed replacing the trunk network with POD (PCA) bases. Recently, transformer architectures have been employed to design neural operators, offering a promising direction for operator learning, e.g., (Cao,, 2021; Li et al.,, 2022; Hao et al.,, 2023).
@@
Physics-Informed Neural Networks (PINNs) (Raissi et al.,, 2019) mark a significant advancement in scientific machine learning. PINNs integrate physical laws directly into the learning process, making them effective for solving differential equations and understanding complex physical systems. This methodology is particularly beneficial in scenarios where data is limited or expensive to obtain. Pioneering the concept of PINO, Li et al., (2021) introduced a dual-resolution approach that combines low-resolution empirical data with high-resolution PDE constraints. This method achieves precise emulation of solution operators across various PDE classes. In parallel, physics-informed DONet by Wang et al., (2021) incorporate regularization strategies enforcing physical law adherence into the training of DONets. Zanardi et al., (2023) presented an approach using PINO for simulations in non-equilibrium reacting flows.  Lee et al., (2023) proposed opPINN, a framework combining physics-informed neural networks with operator learning for solving the Fokker-Planck-Landau (FPL) equation.  Rosofsky et al., (2023) provided a review of applications of physics-informed neural operators. However, existing methods demand one should know the physics laws beforehand, which might not be feasible in many practical applications or complex systems. Our method offers a simple and effective framework, enabling the extraction of implicit physics laws directly from data, even when the data is scarce. Empirically, these pseudo physics laws have proven to be highly beneficial in enhancing the performance of operator learning, as demonstrated in Section 5. Many methods have been developed specifically for discovering differential equations from data, including SINDy (Brunton et al.,, 2016) and its variants (Schaeffer,, 2017; Zhang and Ma,, 2020; Lagergren et al.,, 2020), PINN-SR (Chen et al.,, 2021), Bayesian spline learning (Sun et al.,, 2022), and kernel-based equation discovery (Long et al.,, 2024). To ensure interpretability, these approaches typically assume a specific equation form and perform sparse selection from a set of candidate operators. In contrast to these methods, which prioritize interpretability, our approach focuses on enhancing the prediction accuracy of operator learning under limited data. To this end, we utilize a black-box neural network to represent PDEs. While this offers greater flexibility, it comes at the cost of reduced interpretability. Our method employs an alternating update strategy to jointly refine the PDE representation and improve operator learning.
@@
Our work is also related to the cycle consistence framework (Zhu et al.,, 2017) for image-to-image translation. A critical difference is that cycle-consistence performs unpaired image-to-image translation, while our method aims for accurate paired translation (mapping). In cycle-consistence, the translation is viewed successfully as long as the translated images follow the target distribution. Hence, cycle-consistence has a much more relaxed objective. Another key difference is that our method aims to improve the learning of a function-to-function mapping with very limited data— that motivates us to learn a “pseudo” physics representation. The cycle-consistence framework relies on adversarial training which typically requires a large amount of data to obtain successful learning outcomes.
@@ 5 Experiments
We tested on five commonly used benchmark operator learning problems in literature (Li et al., 2020b, ; Lu et al.,, 2022), including Darcy Flow, Nonlinear Diffusion, Eikonal, Poisson and Advection. In addition, we examined our method in an application of fatigue modeling. The task is to predict the stress intensity factor (SIF) for semi-elliptical surface cracks on plates, given three geometric parameters that characterize the cracks (Merrell et al.,, 2024); see Appendix Fig. 4. The SIF plays a critical role in modeling crack growth by quantifying the stress state near the tip of a crack, and hence SIF computation and analysis is extremely important in fatigue modeling and fracture mechanics (Anderson and Anderson,, 2005). The SIF computation is expensive, because it typically needs to run finite element method (FEM) or extended FEM with very fine meshes (Kuna,, 2013). Due to the complex sequence of computational steps involved in SIF calculation, there is no holistic PDE that directly models the relationship between the geometric features and the SIF function. Instead, SIF computation typically relies on numerical methods and the extraction of local stress fields near the crack tip. The details about all the datasets are given in Section A of the Appendix.
@@
For each operator learning benchmark, we simulated 100 examples for testing, and varied the number of training examples from {5, 10, 20, 30}, except for Advection, we ran with {20, 30, 50, 80} training examples. For SIF prediction, which is much more challenging, we experimented with training size from {400, 500, 600}, and employed 200 test examples. We repeated the evaluation for five times, each time we randomly sampled a different training set. The input to the pseudo physics networks $\phi$ includes all the partial derivatives up to the 2nd order. The details about the hyperparameter settings are given in Appendix Section A.7.
@@
5.1 Results and Analysis
(a) PPI-FNO: learning
(b) PPI-DONet: learning
(c) PPI-FNO: $\lambda$
(d) PPI-DONet: $\lambda$
Figure 3: Learning curve of PPI-FNO on Darcy Flow (a), and of PPI-DONet on nonlinear diffusion (b). In (c) and (d) we show how the weight $\lambda$ of “pseudo physics” affects the operator learning performance. The horizontal line in (c) and (d) are the relative $L_{2}$ errors of standard FNO and DONet.
@@ 5.1 Results and Analysis
Predictive performance with limited data. We reported the average relative $L_{2}$ error and the standard deviation (with and without using pseudo physics informed learning) in Table 1 and Table 2. The model trained with the pseudo physics network (see (5)) is denoted as PPI-FNO or PPI-DONet, short for Pseudo Physics Informed FNO/DONet. In all the cases, with our pseudo physics informed approach, the prediction error of both FNO and DONet experiences a large reduction. For instance, across all training sizes in Darcy Flow and nonlinear diffusion, PPI-FNO reduces the relative $L_{2}$ error of the ordinary FNO by over 60% and 93%, respectively. In Darcy Flow with training sizes 10 to 30, PPI-DONet reduces the error of the ordinary DONet by over 50%. In SIF prediction, our method applied to both FNO and DONet reduced the error by over 30% with training size 400 and 500. Even the minimum reduction across all cases achieves 14.12% (PPI-DONet over DONet on Poisson with training size 5).
Together these results demonstrate the strong positive impact of the learned physics by our neural network model $\phi$ specified in Section 3.1. Although it remains opaque and non-interpretable, it encapsulates valuable knowledge that greatly enhances the performance of operator learning with limited data.
Next, we assessed the accuracy of the learned physics laws by examining the relative $L_{2}$ error in predicting the source functions $f$ from $\phi$ (see (3)). We tested on Darcy Flow, nonlinear diffusion, and Eikonal. We compared a baseline method that removes the convolution layer of $\phi$, leaving only the fully connected layers, namely MLP (see Fig 1 bottom). The results are reported in Appendix Table 5. It can be observed that in nearly every case, adding a convolution layer indeed significantly improves the accuracy of $\phi$. This improvement might be attributed to the convolution layer’s ability to integrate neighboring information and compensate for the information loss introduced by finite difference in approximating the derivatives. We also experimented with multiple convolution layers, but the improvement was found to be marginal.
In addition, we found the operator learning improvement is relatively robust to the accuracy of our physics representation $\phi$. For instance, on Darcy Flow with training size 5 and 10, the relative $L_{2}$ error of $\phi$ network is 0.2285 and 0.1392, which is significantly bigger than with training size 30 where the relative $L_{2}$ error is 0.0688. Yet the error reduction upon FNO (see Table 3(a)) under all the three training sizes is above 60%. The error reduction upon DONet is 40% for training size 5 and over 50% for training size 10 and 30. The results imply that even roughly capturing the underlying physics (with $\phi$) can substantially boost the operator learning performance. In Appendix Section B, we we provide a detailed comparison of the point-wise error of our method against standard NO, along with additional discussions.
Training with rich data. Although our method is designed to enhance operator learning with limited data, we also evaluated its performance when the training data is abundant. Specifically, we increased the training size of the five operator learning benchmarks to 600 and 1000 examples. The results, presented in Appendix Table 3, show that the accuracy of the standard NO and our method, PPI-NO, becomes comparable. For Darcy flow and Advection, our method achieves slight improvement, whereas for Eikonal and Nonlinear diffusion, it performs slightly worse. This may be because, with sufficiently rich data, the standard NO can already utilize the data effectively to achieve good performance, leaving little room for the added value of the learned physics to further improve the results. In such cases, the benefit of incorporating physics knowledge becomes marginal. Furthermore, our alternating updating mechanism brings additional optimization workload, which may introduce additional complexity to the learning process.
Comparison with NO training using ground-truth physics. Our work is motivated by scenarios where the underlying physics is unknown. It is therefore insightful to compare our approach with training NO using the ground-truth physics, as done in PINO. This comparison allows us to examine how our learned “pseudo” physics differs from the true physics in facilitating operator learning. To this end, we tested PINO on the Poisson and Advection cases, using finite difference approximations to compute the derivatives. We used FNO as the baseline. The PDE residual was incorporated into the training loss. As shown in Appendix Table 4, while PINO consistently achieves even better results, the performance of our method is close. This demonstrates that our learned black-box representation, despite lacking interpretability, can provide a similar benefit in improving operator learning performance.
Learning behavior. We examined the learning behavior of our method, which conducts an iterative, alternatingly fine-tuning process. We employed one Darcy Flow, one nonlinear diffusion and one Eikonal dataset, each with 30 examples. We show the test relative $L_{2}$ error along with the iterations in Fig. 3(a), 3(b), and Appendix Fig. 6(a) and Fig. 6(b). As we can see, the predictive performance of our algorithm kept improving and tended to converge at last, affirming the efficacy of the learning process.
Ablation study on pseudo physics network $\phi$. To confirm the efficacy of our designed network $\phi$ in facilitating operator learning, we considered alternative designs for $\phi$: (1) using standard FNO to predict $f$ directly from $u$; no derivative information is included in the input; (2) removing the convolution layer in our model, and just keeping the fully-connected layers, namely MLP; the derivative information is not included in the input. With different designs of $\phi$, we evaluated the PPI learning performance on the Darcy Flow benchmark. The relative $L_{2}$ error in predicting $f$ via $\phi$ and predicting $u$ is reported in Appendix Table 6. It can be seen that our design of $\phi$ consistently outperforms alternative architectures by a notable margin, showing the effectiveness of learning a (black-box) PDE representation and improving the operator learning.
Ablation study on the choice of derivatives. We further investigated the PPI learning performance with respect to the choice of derivatives used in our pseudo physics network. Specifically, we tested PPI-FNO on the Darcy-flow benchmark and varied the order of derivatives up to 0, 1, 2, and 3. The performance is reported in Appendix Table 7. We can see that although the accuracy of $\phi$ with derivatives up to the third order is slightly better than with derivatives up to the second order, the best operator learning performance was still achieved using derivatives up to the second order (which was used in our evaluations). This might be because higher-order derivative information can cause overfitting in the pseudo physics network $\phi$ to a certain degree. Such higher-order information may not be critical to the actual mechanism of the physical system and can therefore impede the improvement of operator learning performance.
Ablation study on the weight $\lambda$. We examined the effect of the weight $\lambda$ of our “pseudo physics”; see (5). To this end, we used one Darcy Flow dataset and nonlinear diffusion dataset with training size 30. We varied $\lambda$ from $[0.5,10^{2}]$, and run PPI-FNO and PPI-DONet on Darcy Flow and nonlinear diffusion, respectively. As shown in Fig. 3(c) and 3(d), we can see that across a wide range of $\lambda$ values, PPI-FNO and PPI-DONet can consistently outperform the standard FNO and DONet respectively by a large margin. However, the choice $\lambda$ does have a significant influence on the operator learning performance, and the best choice is often in between. In Appendix Fig. 6(c) and 6(d), we show results on Ekonal, which we make similar observations.
Memory Cost. Our PPI-NO is more costly than standard NO since we need to train an additional “pseudo” physics network $\phi$ along with the NO model. However, the network $\phi$ is small as compared to the NO component — $\phi$ is simply a pixel-wise MLP coupled with one convolution filter, resulting in a marginal increase in memory cost. Appendix Table 8 shows the parameter count of FNO, DONet and their pseudo-physics-informed version. On average, PPI-FNO increases the number of parameters over FNO by 1.29% while PPI-DONet over DONet by 1.89%.
6 Conclusion
7 Impact Statements
This paper presents work whose goal is to advance machine learning for operator learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix
Appendix A Experimental Details
A.1 Darcy Flow
We considered a steady-state 2D Darcy Flow equation (Li et al., 2020b, ),
$-\nabla\cdot(a(x)\nabla u(x))=f(x)\;\;x\in(0,1)^{2},$
$u(x)=0\;\;x\in\partial(0,1)^{2},$ (6)
where $u({\bf x})$ is the velocity of the flow, $a({\bf x})$ characterizes the conductivity of the media, and $f({\bf x})$ is the source function that can represent flow sources or sinks within the domain. In the experiment, our goal is to predict the solution $u$ given the external source $f$. To this end, we fixed the conductivity $a$, which is generated by first sampling a Gauss random field $\alpha$ in the domain and then applying a thresholding rule: $a({\bf x})=4$ if $\alpha({\bf x})<0$, otherwise $a({\bf x})=12$. We then used another Gauss random field to generate samples of $f$. We followed (Li et al., 2020b, ) to solve the PDE using a second-order finite difference solver and collected the source and solution at a $128\times 128$ grid.
A.2 Nonlinear Diffusion PDE
We next considered a nonlinear diffusion PDE,
$u(-1,t)$ $=u(1,t)=0,\;\;u(x,0)=0,$ (7)
where $(x,t)\in[-1,1]\times[0,1]$. Our objective is to predict the solution function $u$ given the source function $f$. We used the solver provided in (Lu et al.,, 2022), and discretized both the input and output functions at a $128\times 128$ grid. The source $f$ was sampled from a Gaussian process with an isotropic square exponential (SE) kernel for which the length scale was set to 0.2.
A.3 Eikonal Equation
$|\nabla u({\bf x})|=\frac{1}{f({\bf x})},\;\;{\bf x}\in[0,256]\times[0,256],$ (8)
where $u({\bf x})$ is the travel time of the wavefront from the source to location ${\bf x}$, ||$\cdot$|| ⋅ | denotes the Euclidean norm, and $f({\bf x})>0$ is the speed of the wave at ${\bf x}$.
$f({\bf x})=\max(g({\bf x}),0)+1.0,$
where $g(\cdot)$ is sampled from a Gaussian process using the isotropic SE kernel with length-scale $0.1$. We employed the eikonalfm library (https://github.com/kevinganster/eikonalfm/tree/master) that implements the Fast Marching method (Sethian,, 1999) to compute the solution $u$.
A.4 Poisson Equation
Fourth, we considered a 2D Poisson Equation,
A.5 Advection Equation
Fifth, we considered a wave advection equation,
A.6 Fatigue Modeling
Figure 4: Example of semi-elliptic surface crack on a plates (Merrell et al.,, 2024).
We considered predicting the SIF values along semi-elliptic surface cracks on plates, as shown in Fig 4. The SIF value can be viewed as a function of the angle $\phi\in[0,\pi]$, which decides the location of each point on the crack surface. The geometry parameters that characterize the crack shape and position were used as the input, including $a/c$, $a/t$ and $c/b$. In the operator learning framework, the input can be viewed as a function with three constant outputs. The dataset was produced via a high-fidelity FE model under Mode I tension (Merrell et al.,, 2024). Each data instance includes 128 samples of the SIF values drawn uniformly across the range of $\phi$.
A.7 Hyperparameter Settings
For the pseudo physics neural network $\phi$ — see (3) — we tuned the kernel size from {(3, 3), (5, 5), (7, 7), (9, 9)}. The stride was set to 1 and padding was set to “same” to ensure the output shape does not change. In the subsequent fully connected layers, we chose the number of layers from {3, 4, 5, 6}, and the layer width from {16, 32, 64}. We used GeLU activation. For FNO, we set the number of modes to 12 and channels to 32 (in the lifted space). We varied the number of Fourier layers from {2, 3, 4}. For DONet, in all the cases except Darcy Flow, the trunk net and branch net were constructed as fully connected layers. We varied the number of layers from {2, 3, 4} and the layer width was chosen from {30, 40, 50, 60}, with ReLU activation. For the case of Darcy flow, we found that DONet with only fully connected layers exhibited inferior performance. To address this, we introduced convolution layers into the branch net. We selected the number of convolution layers from {3,5,7}, and employed batch normalization and leaky ReLU after each convolution layer. To incorporate the learned pseudo physics representation into the training of FNO or DONet, we randomly sampled 200 input functions to construct the second loss term in (5). We set the maximum number of iterations to $10$ and selected the weight $\lambda$ from $[10^{-1},10^{2}]$. All the models were implemented by PyTorch (Paszke et al.,, 2019), and optimized with ADAM (Kingma and Ba,, 2014). The learning rate was selected from $\{10^{-4},5\times 10^{-4},10^{-3}\}$. The number of epochs for training or fine-tuning FNO, DONet and pseudo physics network $\phi$ was set to 500 to ensure convergence.
@@ A.7 Hyperparameter Settings
PI-FNO 0.3628 $\pm$ 0.0082 0.3205 $\pm$ 0.0121 0.$2$$\pm$ 0.0074 0.1668 $\pm$ 0.0057
Table 5: Relative $L_{2}$ error of using the learned back-box PDE network (3) to predict the input function $f$.
(a) Training size=10
Benchmark MLP Ours
Darcy Flow 0.1819$\pm$0.0026 0.1392$\pm$ 0.0080
Nonlinear Diffusion 0.0660$\pm$0.0069 0.0233$\pm$0.0005
Eikonal 0.0144$\pm$0.0009 0.0108 $\pm$ 0.0006
(b) Training size=30
Benchmark MLP Ours
Darcy Flow 0.1413$\pm$0.0013 0.0688$\pm$ 0.0032
Nonlinear Diffusion 0.0463$\pm$0.0022 0.0163$\pm$0.0002
Eikonal 0.0070$\pm$0.00005 0.0052 $\pm$ 0.0002
(a) Predicting $f$ via $\phi$ with different architectures.
Training size 5 10 20 30
FNO 0.7229$\pm$0.0318 0.5759$\pm$ 0.0126 0.4257$\pm$ 0.0106 0.3160$\pm$ 0.0037
MLP 0.7169$\pm$0.0160 0.6598$\pm$ 0.0056 0.6464$\pm$ 0.0029 0.6277$\pm$ 0.0032
(b) Predicting $u$.
Training size 5 10 20 30
PPI-FNO with FNO as $\phi$ 0.5853$\pm$0.0153 0.3871$\pm$ 0.0124 0.2613$\pm$ 0.0190 0.1629$\pm$ 0.0064
PPI-FNO with MLP as $\phi$ 0.7262$\pm$0.0920 0.5516$\pm$ 0.0699 0.4568$\pm$ 0.0857 0.3983$\pm$ 0.1051
(a) Predicting $f$ via $\phi$.
Training size 5 10 20 30
order 0 0.7126$\pm$0.0131 0.5733$\pm$0.0208 0.4812$\pm$0.0399 0.3445$\pm$0.0182
order $\leq$ 1 0.2926$\pm$0.0118 0.2006$\pm$0.0047 0.1379$\pm$0.0051 0.1084$\pm$0.0053
order $\leq 2$ 0.2285$\pm$0.0147 0.1392$\pm$0.0080 0.0898$\pm$0.0046 0.0688$\pm$0.0032
order $\leq 3$ 0.2058$\pm$0.0192 0.1123$\pm$0.0039 0.0712$\pm$0.0021 0.0585$\pm$0.0030
(b) Predicting $u$.
Training size 5 10 20 30
order 0 0.6352$\pm$0.0673 0.4523$\pm$0.0621 0.3570$\pm$0.0658 0.2737$\pm$0.0643
order $\leq$ 1 0.3386$\pm$0.0259 0.2161$\pm$0.0083 0.1645$\pm$0.0114 0.1197$\pm$0.0132
order $\leq 2$ 0.1716$\pm$0.0048 0.0956$\pm$0.0084 0.0680$\pm$0.0031 0.0642$\pm$0.0010
order $\leq 3$ 0.2959$\pm$0.0381 0.1719$\pm$0.0213 0.1193$\pm$0.0158 0.0828$\pm$0.0054
Parameter count FNO PPI-FNO (increase) DONet PPI-DONet (increase)
Eikonal 1,188,353 1,197,220 (+0.75%) 824,501 833,368 (+1.08%)
Poisson 1,188,353 1,197,220 (+0.75%) 824,501 833,368 (+1.08%)
Advection 1,188,353 1,197,220 (+0.75%) 210,101 218,968 (+4.22%)
Appendix B Pointwise Error
@@
For a further assessment, we conducted a fine-grained evaluation by visualizing the predictions and point-wise errors made by each method. In Fig. 7(a) and 7(b), we showcased the predictions and point-wise errors using PPI-DONet for Darcy Flow, PPI-FNO for nonlinear diffusion, respectively. Additional examples of predictions and point-wise errors are provided in Fig. 5, 8(a), 8(b), 9(a), and 9(b).
@@
It is evident that without the assistance of the pseudo physics laws learned by our method, the ordinary DONet and FNO frequently missed crucial local structures, sometimes even learning entirely incorrect structures. For example, In Fig. 7(a) the first row, DONet missed one mode, while in the second and third row of Fig. 7(a), DONet failed to capture all the local modes. After incorporating the learned physics, DONet (now denoted as PPI-DONet; see the third column) successfully captures all the local modes, including their shapes and positions. Although not all the details are exactly recovered, the point-wise error is substantially reduced, particularly in those high error regions of the ordinary DONet; see the fourth column of Fig. 7(a). In another instance, as shown in Fig. 7(b), where the ordinary FNO (second column) captured the global shape of the solution, but the mis-specification of many local details led to large point-wise errors across many regions (fourth column). In contrast, PPI-FNO (third column) not only identified the structures within the solution but also successfully recovered the details. As a result, the point-wise error (fifth column) was close to zero everywhere. Additional instances can be found in Fig. 8(a), the first three rows illustrate that ordinary FNO (trained with 5, 10, and 20 examples, respectively) estimates an entirely incorrect structure of the solution, indicating that the training data is insufficient for FNO to capture even the basic structure of the solution. In contrast, after fine-tuning with our learned physics laws from the same sparse data, PPI-FNO accurately figured out the solution structures and yielded a substantial reduction in point-wise error across nearly everywhere. The point-wise error became uniformly close to zero. With 30 examples, the ordinary FNO was then able to capture the global structure of the solution, but the details in the bottom left, bottom right, and top right corners were incorrectly predicted. In comparison, PPI-FNO further recovered these details accurately.
(a) PPI-FNO: learning
(b) PPI-DONet: learning
(c) PPI-FNO: $\lambda$
(d) PPI-DONet: $\lambda$
Figure 6: Learning curve of PPI-FNO (a) and PPI-DONet (b) on Eikonal with 30 training examples. Shown in (c) and (d) is how the weight $\lambda$ of “pseudo physics” affects the operator learning performance. The horizontal lines in (c) and (d) are the relative $L_{2}$ error of standard FNO and DONet, respectively.

### v2

@@ 4 Related Work
Operator learning is a rapidly evolving research field, with a variety of methods categorized as neural operators. Alongside FNO, several notable approaches have been introduced, such as low-rank neural operator (LNO) (Li et al., 2020b), multipole graph neural operator (MGNO) (Li et al., 2020a), multiwavelet-based NO (Gupta et al., 2021), and convolutional NOs (CNO) (Raonic et al., 2023). Deep Operator Net (DON) (Lu et al., 2021) is another popular approach, consisting of a branch network applied to input function values and a trunk network applied to sampling locations. The final prediction is obtained through the dot product of the outputs from the two networks. To enhance stability and efficiency, Lu et al. (2022) proposed replacing the trunk network with POD (PCA) bases. Recently, transformer architectures have also been employed to design neural operators, e.g., (Cao, 2021; Li et al., 2022; Hao et al., 2023).
@@
Physics-Informed Neural Networks (PINNs) (Raissi et al., 2019) mark a significant advancement in scientific machine learning. PINNs integrate physical laws directly into the learning process, making them effective for solving differential equations and understanding complex physical systems. This methodology is particularly beneficial in scenarios where data is limited or expensive to obtain. Li et al. (2021) introduced a dual-resolution approach that combines low-resolution empirical data with high-resolution PDE constraints. This method achieves precise emulation of solution operators across various PDE classes. In parallel, physics-informed DONet by Wang et al. (2021) incorporate regularization strategies enforcing physical law adherence into the training of DONets. Zanardi et al. (2023) presented an approach using PINO for simulations in non-equilibrium reacting flows.  Lee et al. (2023) proposed opPINN, a framework combining physics-informed neural networks with operator learning for solving the Fokker-Planck-Landau (FPL) equation.  Rosofsky et al. (2023) provided a review of applications of physics-informed neural operators. However, existing methods demand one must know the physics laws beforehand, which might not be feasible in many practical applications or complex systems. Our method offers a simple and effective framework, enabling the extraction of implicit physics laws directly from data, even when the data is scarce. Empirically, these pseudo physics laws have proven to be highly beneficial in enhancing the performance of operator learning, as demonstrated in Section 5. Many methods have been developed specifically for discovering differential equations from data, including SINDy (Brunton et al., 2016) and its variants (Schaeffer, 2017; Zhang & Ma, 2020; Lagergren et al., 2020), PINN-SR (Chen et al., 2021), Bayesian spline learning (Sun et al., 2022), and kernel-based equation discovery (Long et al., 2024). To ensure interpretability, these approaches typically assume a specific equation form and perform sparse selection from a set of candidate operators. In contrast to these methods, which prioritize interpretability, our approach focuses on enhancing the prediction accuracy of operator learning under limited data. To this end, we utilize a black-box neural network to represent PDEs. While this offers greater flexibility, it comes at the cost of reduced interpretability. Our method employs an alternating update strategy to jointly refine the PDE representation and improve operator learning.
@@
Our work is also related to the cycle consistence framework (Zhu et al., 2017) for image-to-image translation. A critical difference is that cycle-consistence performs unpaired image-to-image translation, while our method aims for accurate paired translation (mapping). In cycle-consistence, the translation is viewed successfully as long as the translated images follow the target distribution. Hence, cycle-consistence has a much more relaxed objective. Another key difference is that our method aims to improve the learning of a function-to-function mapping with very limited data— that motivates us to learn a “pseudo” physics representation. The cycle-consistence framework relies on adversarial training which typically requires a large amount of data to obtain successful learning outcomes.
@@ 5 Experiments
We tested on five commonly used benchmark operator learning problems in the literature (Li et al., 2020b; Lu et al., 2022), including Darcy Flow, Nonlinear Diffusion, Eikonal, Poisson and Advection. In addition, we examined our method in an application of fatigue modeling. The task is to predict the stress intensity factor (SIF) for semi-elliptical surface cracks on plates, given three geometric parameters that characterize the cracks (Merrell et al., 2024); see Fig. 3. The SIF plays a critical role in modeling crack growth by quantifying the stress state near the tip of a crack, and hence SIF computation and analysis are extremely important in fatigue modeling and fracture mechanics (Anderson & Anderson, 2005). The SIF computation is expensive, because it typically needs to run finite element method (FEM) or extended FEM with very fine meshes (Kuna, 2013). Due to the complex sequence of computational steps involved in SIF calculation, there is no holistic PDE that directly models the relationship between the geometric features and the SIF function. Instead, SIF computation typically relies on numerical methods and the extraction of local stress fields near the crack tip. The details about all the datasets are given in Section A of the Appendix.
Figure 3: Example of semi-elliptic surface crack on a plates (Merrell et al., 2024).
@@
For each operator learning benchmark, we simulated 100 examples for testing, and varied the number of training examples from {5, 10, 20, 30}, except for Advection, we ran with {20, 30, 50, 80} training examples. For SIF prediction, which is much more challenging, we experimented with training size from {400, 500, 600}, and employed 200 test examples. We repeated the evaluation for five times, each time we randomly sampled a different training set. The input to the pseudo physics networks $\phi$ includes all the partial derivatives up to the 2nd order. For the pseudo physics neural network $\phi$ — see equation 3 — we tuned the kernel size from {(3, 3), (5, 5), (7, 7), (9, 9)}. The stride was set to 1 and padding was set to “same” to ensure the output shape does not change. In the subsequent fully connected layers, we chose the number of layers from {3, 4, 5, 6}, and the layer width from {16, 32, 64}. We used GeLU activation. For FNO, we set the number of modes to 12 and channels to 32 (in the lifted space). We varied the number of Fourier layers from {2, 3, 4}. For DONet, in all the cases except Darcy Flow, the trunk net and branch net were constructed as fully connected layers. We varied the number of layers from {2, 3, 4} and the layer width was chosen from {30, 40, 50, 60}, with ReLU activation. For the case of Darcy flow, we found that DONet with only fully connected layers exhibited inferior performance. To address this, we introduced convolution layers into the branch net. We selected the number of convolution layers from {3,5,7}, and employed batch normalization and leaky ReLU after each convolution layer. To incorporate the learned pseudo physics representation into the training of FNO or DONet, we randomly sampled 200 input functions to construct the second loss term in equation 5. We set the maximum number of iterations to $10$ and selected the weight $\lambda$ from $10^{-1},10^{2}]$. All the models were implemented by PyTorch (Paszke et al., 2019), and optimized with ADAM (Kingma & Ba, 2014). The learning rate was selected from $\{10^{-4},5\times 10^{-4},10^{-3}\}$. The number of epochs for training or fine-tuning FNO, DONet and pseudo physics network $\phi$ was set to 500 to ensure convergence.
@@
5.1 Predictive Performance
@@
Table 3: Relative $L_{2}$ error of using the learned back-box PDE network equation 3 to predict the input function $f$.
(a) Training size=10
Benchmark MLP Ours
[…]
(b) Training size=30
Benchmark MLP Ours
[…]
@@
PI-FNO 0.3628 $\pm$ 0.0082 0.3205 $\pm$ 0.0121 0.2222 $\pm$ 0.0074 0.1668 $\pm$ 0.0057
We reported the average relative $L_{2}$ error and the standard deviation (with and without using pseudo physics informed learning) in Table 1 and Table 2. The model trained with the pseudo physics network (see equation 5) is denoted as PPI-FNO or PPI-DONet, short for Pseudo Physics Informed FNO/DONet. As shown, across all the cases, with our pseudo physics informed approach, the prediction error of both FNO and DONet undergoes a large reduction. For instance, across all training sizes in Darcy Flow and nonlinear diffusion, PPI-FNO reduces the relative $L_{2}$ error of the ordinary FNO by over 60% and 93%, respectively. In Darcy Flow with training sizes 10 to 30, PPI-DONet reduces the error of the ordinary DONet by over 50%. In SIF prediction, our method applied to both FNO and DONet reduced the error by over 30% for training sizes of 400 and 500. Even the minimum reduction across all cases achieves 14.12% (PPI-DONet over DONet on Poisson with training size 5).
Together these results demonstrate the strong positive impact of the learned physics by our neural network model $\phi$ specified in Section 3.1. Although it remains opaque and non-interpretable, it encapsulates valuable knowledge that substantially enhances the performance of operator learning with limited data.
Next, we assessed the accuracy of the learned physics laws by examining the relative $L_{2}$ error in predicting the source functions $f$ from $\phi$ (see equation 3). We tested on Darcy Flow, nonlinear diffusion, and Eikonal. We compared a baseline method that removes the convolution layer of $\phi$, leaving only the fully connected layers, namely MLP (see Fig 1 bottom). The results are reported in Table 3. It can be observed that in nearly every case, adding a convolution layer indeed significantly improves the accuracy of $\phi$. This improvement might be attributed to the convolution layer’s ability to integrate neighboring information and compensate for the information loss introduced by finite difference in approximating the derivatives. We also experimented with multiple convolution layers, but the improvement was found to be marginal.
In addition, we found the operator learning improvement is relatively robust to the accuracy of our physics representation $\phi$. For instance, on Darcy Flow with training size 5 and 10, the relative $L_{2}$ error of $\phi$ network is 0.2285 and 0.1392, which is significantly bigger than with training size 30 where the relative $L_{2}$ error is 0.0688. Yet the error reduction upon FNO (see Table 1(a)) under all the three training sizes is above 60%. The error reduction upon DONet is 40% for training size 5 and over 50% for training size 10 and 30. The results imply that even roughly capturing the underlying physics (with $\phi$) can substantially boost the operator learning performance.
@@
For a further assessment, we conducted a fine-grained evaluation by visualizing the predictions and point-wise errors made by each method. In Fig. 4(a) and 4(b), we showcased the predictions and point-wise errors using PPI-DONet for Darcy Flow, PPI-FNO for nonlinear diffusion, respectively. Additional examples of predictions and point-wise errors are provided in Fig. 5(a), 5(b), 6(a), 6(b), and 7.
@@
It is evident that without the assistance of the pseudo physics laws learned by our method, the ordinary DONet and FNO frequently missed crucial local structures, sometimes even learning entirely incorrect structures. For example, In Fig. 4(a) the first row, DONet missed one mode, while in the second and third row of Fig. 4(a), DONet failed to capture all the local modes. After incorporating the learned physics, DONet (now denoted as PPI-DONet; see the third column) successfully captures all the local modes, including their shapes and positions. Although not all the details are exactly recovered, the point-wise error is substantially reduced, particularly in those high error regions of the ordinary DONet; see the fourth column of Fig. 4(a). In another instance, as shown in Fig. 4(b), where the ordinary FNO (second column) captured the global shape of the solution, but the mis-specification of many local details led to large point-wise errors across many regions (fourth column). In contrast, PPI-FNO (third column) not only identified the structures within the solution but also successfully recovered the details. As a result, the point-wise error (fifth column) was close to zero everywhere. Additional instances can be found in Fig. 5(a), the first three rows illustrate that ordinary FNO (trained with 5, 10, and 20 examples, respectively) estimates an entirely incorrect structure of the solution, indicating that the training data is insufficient for FNO to capture even the basic structure of the solution. In contrast, after joint training with our physics network from the same sparse data, PPI-FNO accurately figured out the solution structures and yielded a substantial reduction in point-wise error across nearly everywhere. The point-wise error became uniformly close to zero. With 30 examples, the ordinary FNO was then able to capture the global structure of the solution, but the details in the bottom left, bottom right, and top right corners were incorrectly predicted. In comparison, PPI-FNO further recovered these details accurately.